f(x) = 3x – 7 and g(x) = –2x – 6. Find (ƒ ∘ g)(4)
Answer by santukeerti(5)

(fo g)(4)
f(g(4))
but g(x)=-2x-6
so g(4)=-2*(4)-6
g(4)=-8-6
g(4)=-14
now
f(g(4))=f(-14)
but f(x)=3x-7
so f(-14)=3*(-14)-7
f(-14)=-42-7
f(-14)=-49
therefore (f o g)(4)=-49